Michelle is 7 years older than her sister Joan, and Joan is 3 years younger than their brother Ryan. If the sum of their ages is 64, how old is Joan?
Answer by algebra hello(55)

let ryan's age =x
then joan's age =x-3
and Michelle's age=x-3+7=x+4
now according to question,
x+(x-3)+(x+4)=64
3x+1=64
3x=63
x=21
Joan's age=21-3=18